# Uncertainty-Aware Collaborative System of Large and Small Models for Multimodal Sentiment Analysis

### v1

@@ Introduction
where $Fusion(\cdot)$ is a function that integrates the concatenated features into a fused representation of dimension $d_{m}$. The fused representation $\bm{h}_{m}$ is then fed into a prediction function $Predictor(\cdot)$ to produce the final sentiment score.
Uncertainty Estimation
@@
(1) Prediction-Truth Differences, which employs an additional predictor to measure the absolute difference between predicted and true labels, treating larger differences as indicators of higher uncertaint. However, predicting label differences is very difficult, requiring the model to have full discriminative power to estimate the labels of all samples.
(2) Ensemble Variance, which trains multiple diverse predictors on different data subsets and uses the variance among their predictions as an uncertainty measure. Although this approach is theoretically sound, it introduces computational overhead because each input must be processed by multiple predictors, conflicting with our efficiency goals.
(3) Our Method – Classification-based Entropy:
To overcome these limitations, we propose a classification-based entropy approach. Specifically, we transform the regression task into a three-class classification problem and employ a classifier to conduct the classification task, enabling the estimation of predictive uncertainty by computing the entropy of the output probability distribution. This method offers several key advantages: (i) computational efficiency achieved through a single forward pass; (ii) a principled, information-theoretic interpretation of uncertainty; and (iii) reducing the difficulty in uncertainty estimation, as the uncertainty of classification tasks is easier to estimate.
Specifically, alongside the primary regression predictor that outputs continuous sentiment scores $\hat{y}_{s}$, we introduce an auxiliary classifier. The continuous sentiment scores are discretized into three categories according to their polarity:
$class=\begin{cases}1\;(Negative),&y<0\\ 0\;(Neutral),&y=0\\ 2\;(Positive),&y>0\end{cases}$ (4)
@@
$\begin{split}\bm{z}_{cls}&=Classifier(\bm{h}_{m}),\\ p_{s}&=Softmax(\bm{z}_{cls})=[p_{1},p_{2},p_{3}]\end{split}$ (5)
where $p_{s}$ denotes the predicted probability distribution over the three sentiment classes. The model is jointly optimized with a combined loss that integrates both regression and classification objectives:
@@
where $\mathcal{L}_{reg}$ represents the Mean Squared Error (MSE) for regression, $\mathcal{L}_{cls}$ is the Cross-Entropy (CE) loss for classification, and $y_{cls}$ is the discretized ground truth label. During inference, uncertainty is quantified by the entropy of the predicted class distribution:
$u_{s}=H(p_{s})=-\sum_{i=1}^{3}p_{s,i}\cdot\log(p_{s,i})$ (7)
where $u_{s}$ denotes the uncertainty score, $H(\cdot)$ is the entropy function, and $p_{s,i}$ is the predicted probability of the $i$-th sentiment class. This entropy-based metric naturally reflects the model’s confidence: a high entropy value indicates greater uncertainty, corresponding to a more uniform distribution over classes, while a low entropy implies strong confidence concentrated on a single class (Xue et al. 2025).
MLLM Refinement
We utilize HumanOmni (Yang, Dong, and Qiang 2025), a 7-billion-parameter MLLM, as the expert oracle for handling difficult samples. The model’s output is formulated as:
$\hat{y}_{l}=\text{HumanOmni}(\bm{U}_{t},\bm{U}_{a},\bm{U}_{v})$ (8)
where $\bm{U}_{t},\bm{U}_{a},\bm{U}_{v}$ denote the input features from text, audio, and visual modalities, respectively, and $\hat{y}_{l}$ is the predicted sentiment score. To efficiently adapt HumanOmni to our task, we apply Low-Rank Adaptation (LoRA) (Hu et al. 2022), which decomposes the weight update into low-rank matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times d}$, with rank $r$ much smaller than $d$. The adapted weight matrix is given by $W_{new}=W+AB$, where $W$ is the original weight matrix and only $A$ and $B$ are trainable during fine-tuning. The MLLM then outputs a probability distribution over the possible sentiment classes, computed as:
@@
$\mathbf{p}_{l}=\mathrm{Softmax}(\mathrm{logits}_{l})=[p_{1},p_{2},\ldots,p_{n}],$ (9)
@@
where $\mathbf{p}_{l}$ represents the predicted probability vector, $p_{i}$ is the probability of the $i$-th class, and $n$ is the total number of sentiment classes. We quantify uncertainty by calculating the entropy of this distribution:
$u_{l}=H(\mathbf{p}_{l})=-\sum_{i=1}^{n}p_{l,i}\cdot\log(p_{l,i}),$ (10)
@@
where $u_{l}$ denotes the uncertainty score and $H(\cdot)$ is the entropy function. Higher entropy corresponds to higher uncertainty, providing a principled confidence measure for the model’s predictions. As output consists of multiple tokens, we average the uncertainties of all label-related output tokens to obtain the final uncertainty.
Uncertainty Threshold Estimation
@@
We adopt a principled statistical approach to determine the uncertainty thresholds for both the small model and the MLLM. This method analyzes the uncertainty distributions on the validation set, providing a theoretically grounded yet computationally efficient way to select thresholds.
@@
The threshold selection procedure is applied consistently to both models. For each model, we partition the validation samples into two groups based on whether the model’s predicted polarity matches the ground truth. We define $\mathbb{D}_{same}$ as the set of samples where the prediction and ground truth share the same polarity (both positive or both negative), and $\mathbb{D}_{opposite}$ as the set where they differ in polarity. For each group, we fit a Gaussian distribution to the uncertainty values and compute the corresponding mean uncertainties:
@@
where $(u)_{i}$ denotes the uncertainty for sample $i$, and the means are estimated via Gaussian fitting on the validation set. Finally, we apply this method with Gaussian fitting to compute the specific thresholds for both models:
@@
Summary of U-ACS
Building upon the uncertainty thresholds $\tau_{1}$ and $\tau_{2}$, we now present the overall inference workflow, which is shown in Algorithm 1. This workflow employs a dynamic collaboration strategy that sequentially leverages UBM and MLLM, guided by their uncertainty estimates, to efficiently and accurately produce the final prediction. The process proceeds through three stages:
Stage 1: Small Model Filtering. Compute the baseline prediction $\hat{y}_{s}$ and uncertainty $u_{s}$. If the uncertainty is below the adaptive validation-calibrated threshold $\tau_{1}$, accept this prediction directly. Otherwise, proceed to the MLLM Refinement Stage:
$\hat{y}=\begin{cases}\hat{y}_{s},&u_{s}\leq\tau_{1}\\ \text{Proceed to Stage 2},&u_{s}>\tau_{1}\end{cases}$ (13)
Stage 2: MLLM Refinement. For samples with higher uncertainty, invoke the MLLM to obtain prediction $\hat{y}_{l}$ along with its uncertainty $u_{l}$. Accept the prediction if the uncertainty is below the second threshold $\tau_{2}$:
$\hat{y}=\begin{cases}\hat{y}_{l},&u_{l}\leq\tau_{2}\\ \text{Proceed to Stage 3},&u_{l}>\tau_{2}\end{cases}$ (14)
Stage 3: Collaborative Reasoning. For cases still deemed highly uncertain, apply collaborative strategies based on whether two models agree on sentiment polarity:
  • Agreement (Weighted Averaging): If both models predict the same polarity, compute adaptive weights based on uncertainty and combine their outputs:
    $\begin{split}w_{s}&=\frac{1/(u_{s}+\epsilon)}{1/(u_{s}+\epsilon)+1/(u_{l}+\epsilon)},\\ \hat{y}&=w_{s}\cdot\hat{y}_{s}+(1-w_{s})\cdot\hat{y}_{l}\end{split}$ (15)
  • Disagreement (Cross-Verification): If two models disagree, augment the prompt with previous predictions and uncertainties, and re-predict through the MLLM:
    $\hat{y}=\text{HumanOmni}(X,P_{enhanced})$ (16)
    where $P_{enhanced}$ provides richer context by including UBM’s prediction $\hat{y}_{s}$ and uncertainty $u_{s}$, along with the prediction and uncertainty of the MLLM.
@@
Input: Sample $X=(\bm{U}_{t},\bm{U}_{a},\bm{U}_{v})$, Small model $M_{s}$, MLLM $M_{l}$, uncertainty thresholds $\tau_{1},\tau_{2}$;
@@
11:   if $\hat{y}_{s}$ and $\hat{y}_{l}$ are the same polarity then
@@
20:end if
Experiments
U-ACS has been evaluated using the CMU-MOSI (MOSI) (Zadeh et al. 2016), CMU-MOSEI (MOSEI) (Zadeh et al. 2018), and CH-SIMS v1 (SIMS) (Yu et al. 2020) datasets. Due to space constraints, introductions to the experimental settings, baselines, evaluation metrics, and descriptions of the datasets are provided in the supplementary material.
Experiment Results
To assess the performance of U-ACS, we utilized three well-recognized MSA datasets. As presented in Table 1, U-ACS demonstrates superior performance, achieving state-of-the-art results in some evaluation metrics while maintaining computational efficiency. Compared to HumanOmni, U-ACS achieves 0.5 points improvement in Acc2 and F1 score on MOSI. For MOSEI, U-ACS achieves 0.3 points improvement in Acc2 and 0.4 points improvement in F1 score. Table 2 presents the robustness validation results of U-ACS in cross-lingual scenarios. On the Chinese dataset CH-SIMS v1, the model maintains leading performance across most metrics, including Acc3, Acc2 and F1 score. Furthermore, U-ACS significantly outperforms traditional multimodal methods, demonstrating the effectiveness of uncertainty-guided collaboration across different languages and data characteristics.
Notably, compared with HumanOmni, our method achieves the optimal performance in the Acc2 metric across three datasets. The superiority of the Acc2 performance is attributed to the design of the model in terms of threshold setting. We explicitly divide samples into same and opposite categories according to the coarse-grained positive and negative categories, which is highly consistent with the computational logic of Acc2. Moreover, in the small model UBM, we transform continuous sentiment labels into positive, negative, and neutral classes. These designs make the optimization of Acc2 (positive versus negative) more targeted, thereby highlighting significant advantages.
Method CMU-MOSI CMU-MOSEI
[…]
MISA (Hazarika, Zimmermann, and Poria 2020) 42.3 83.4 83.6 0.783 0.761 52.2 85.5 85.3 0.555 0.756
DBM (Yu et al. 2021) 45.6 84.5 84.4 0.719 0.794 53.7 84.8 84.7 0.535 0.761
[…]
MAG-BERT (Rahman et al. 2020) 43.6 84.4 84.6 0.727 0.781 52.7 84.8 84.7 0.543 0.755
Self-MM (Yu et al. 2021) 45.8 84.9 84.8 0.731 0.785 53.0 85.2 85.2 0.540 0.763
ConFEDE (Yang et al. 2023) 42.3 85.5 85.5 0.742 0.784 54.9 85.8 85.8 0.522 0.780
DLF (Wang et al. 2025a) 47.1 85.1 85.0 0.731 0.781 53.9 85.4 85.3 0.536 0.764
DEVA (Wu et al. 2025) 46.3 86.3 86.3 0.730 0.787 52.3 86.1 86.2 0.541 0.769
HumanOmni (Zhao et al. 2025) 52.8 91.3 91.3 0.549 0.881 58.6 86.1 85.4 0.483 0.807
U-ACS 51.2 91.8 91.8 0.585 0.868 55.4 86.4 85.8 0.506 0.784
Table 1: Performance comparison on CMU-MOSI and CMU-MOSEI datasets. Best results are in bold, second-best are underlined. means the results are from (Feng et al. 2024).
Method CH-SIMS v1
[…]
MISA - - 76.5 76.6 0.447 0.563
DBM 41.5 65.5 80.0 80.4 0.425 0.595
$TFN^{\dagger}$ 39.3 65.1 78.4 78.6 0.432 0.591
$MulT^{\dagger}$ 37.9 64.8 78.6 79.7 0.453 0.564
MAG-BERT - - 74.4 71.8 0.492 0.399
DEVA 43.1 65.4 79.6 80.3 0.424 0.583
$KuDA^{\dagger}$ 43.5 66.5 80.7 80.7 0.408 0.613
HumanOmni 52.1 72.9 85.1 85.0 0.327 0.749
U-ACS 51.0 73.1 85.8 85.1 0.354 0.691
Table 2: Performance comparison on CH-SIMS v1 dataset. Best results are in bold, second-best are underlined. means the results are from (Feng et al. 2024).
Study on Computational Efficiency
Table 3 compares the computational costs, Acc2, and MAE of different approaches. Compared with HumanOmni, our strategy achieves the optimal performance in the Acc2 metric on MOSI, MOSEI, and SIMS datasets, while attaining significant efficiency improvements. Specifically, on the MOSI dataset, the runtime of ours is only 33% of that of the HumanOmni, which significantly reduces the inference time while maintaining the performance. These results indicate that, our strategy, while ensuring that the model performance reaches the level of MLLM, has significantly shortened inference time and reduced model complexity, greatly improving operational efficiency and ultimately achieving a balance between performance and efficiency.
[…]
CMU-MOSI UBM 15.35 86.1 0.716
HumanOmni 379.18 91.3 0.549
Ours 125.76 91.8 0.585
CMU-MOSEI UBM 35.02 86.3 0.527
HumanOmni 9815.56 86.1 0.483
Ours 5561.18 86.4 0.506
CH-SIMS v1 UBM 18.76 80.5 0.409
HumanOmni 470.61 85.1 0.327
Ours 260.58 85.8 0.354
Table 3: Computational efficiency comparison. Time is measured in seconds per sample. Best results are in bold, second-best are underlined.
Study on Uncertainty Distributions
To determine optimal uncertainty thresholds $\tau_{1}$ and $\tau_{2}$, We fit the uncertainty of correctly predicted samples and incorrectly predicted samples to a Gaussian distribution. Figure 3 shows the uncertainty distributions for correct and incorrect predictions on CMU-MOSI testing set. Although there is an overlap in the uncertainties of some samples, the uncertainty distributions of correctly predicted samples and incorrectly predicted samples on the MOSI dataset still show significant differences. The visualization indicates the effectiveness of the proposed uncertainty estimation method (especially for UBM whose uncertainty is hard to estimate). Consequently, effective classification of the two types of samples can be achieved by calculating thresholds.
(a) UBM
(b) MLLM
Figure 3: The distribution of uncertainties on CMU-MOSI.
Case Study
To illustrate the effectiveness of our strategy, we present a case study on a challenging sample from the CH-SIMS dataset. Figure 4 demonstrates how our framework processes a complex sentiment analysis case where individual models struggle but collaboration system succeeds.
Ablation Studies
Dataset Description Acc7 Acc2 MAE
MOSI W/o Cross-Verification 50.4 90.4 0.631
W/o Uncertainty 51.7 89.6 0.600
Prediction-Truth Difference 46.7 87.0 0.649
Ensemble Variance 46.9 88.7 0.632
Ours 51.2 91.8 0.585
Dataset Description Acc5 Acc2 MAE
SIMS W/o Cross-Verification 50.1 85.1 0.358
W/o Uncertainty 47.9 83.6 0.357
Prediction-Truth Difference 45.7 81.2 0.407
Ensemble Variance 41.8 78.1 0.398
Ours 51.0 85.8 0.354
Table 4: Ablation experiments on CMU-MOSI and CH-SIMS. Best results are in bold, second-best are underlined.
We conduct comprehensive ablation studies to validate each component of our framework. Table 4 presents the results:
(3) Uncertainty Quantification Methods: Our entropy-based approach substantially outperforms alternative uncertainty estimation methods. Compared to prediction-truth difference, we achieve 4.8-point improvements in Acc2 on MOSI and 4.6-point improvements on SIMS. Ensemble variance performs even worse, particularly on SIMS with a 7.7-point Acc2 gap and 9.2-point Acc5 gap. These results confirm that our single-pass entropy computation provides both efficiency and effectiveness for uncertainty estimation, which cleverly circumvents the issue of difficulty in computing uncertainty for regression task via transforming regression task into classification task.
Conclusion
We propose a novel method named U-ACS to combine traditional multimodal models with MLLMs to address the imbalance between performance and efficiency in MSA. U-ACS explicitly computes the uncertainty of traditional models and MLLM, designing an uncertainty-driven cascade mechanism to select samples yielding high predictive uncertainty for processing by the MLLM. We also propose advanced collaboration strategies to handle the hardest samples. U-ACS significantly improves efficiency of MLLM while ensuring model performance.
Appendix A Appendix
Datasets
We evaluate our Uncertainty-Aware Collaborative System (U-ACS) on three widely used multimodal sentiment analysis benchmarks: CMU-MOSI, CMU-MOSEI, and CH-SIMS v1. Each dataset presents unique characteristics and challenges for multimodal sentiment analysis.
  • CMU-MOSI (Zadeh et al. 2016) contains 2,199 video segments that study the interaction patterns between facial gestures and spoken words for sentiment prediction. The dataset provides fine-grained sentiment annotations on a continuous scale from strongly negative to strongly positive, making it ideal for evaluating both regression and classification capabilities of multimodal models.
  • CMU-MOSEI (Zadeh et al. 2018) is a larger multimodal dataset for sentiment analysis and emotion recognition, containing 23,453 annotated video segments from 1,000 distinct speakers across 250 topics, serving as the next generation expansion of MOSI. The dataset’s scale and diversity provide a robust evaluation platform for assessing model generalization across different speakers and topics.
  • CH-SIMS v1 (Yu et al. 2020) is a Chinese multimodal dataset which contains 2,281 refined video segments in the wild with both multimodal and independent unimodal annotation. This dataset enables cross-lingual validation of our approach and demonstrates the framework’s robustness across different languages and cultural contexts.
Evaluation Metrics
Following standard multimodal sentiment analysis (MSA) evaluation protocols, we report comprehensive performance metrics across different granularities:
Classification Metrics:

### v2

@@ III-C Pipeline of UBM
where $\bm{x}_{t}$, $\bm{x}_{a}$, $\bm{x}_{v}$ represent unimodal representations (i.e., text, acoustic, visual), $\bm{x}_{concat}$ is the concatenated representation, $Fusion(\cdot)$ is a function that integrates the concatenated features into a fused representation $\bm{h}_{m}$ of dimension $d_{m}$. $W_{1}$ and $b_{1}$ are learnable parameters.
3) Predictor
MSA tasks are generally treated as regression tasks, the sentiment predictor need to predict a sentiment label for each sample, and this label is a continuous value that reflects the fine-grained sentiment intensity of sample:
$\hat{y}_{s}=W_{r}^{(2)}\text{ReLU}(W_{r}^{(1)}\bm{h}_{m}+b_{r}^{(1)})+b_{r}^{(2)}$ (5)
where $\hat{y}_{s}$ is the predicted sentiment label for a single sample; $W_{r}^{(1)}$, $W_{r}^{(2)}$, $b_{r}^{(1)}$, $b_{r}^{(2)}$ are all learnable parameters of the regression predictor, with ReLU acting as the activation function to capture non-linear relationships between multimodal features and sentiment scores. The fused multimodal representation $\bm{h}_{m}$ (in Eq. 4) is fed as input to this predictor.
4) Uncertainty Estimation
@@
  • Prediction-Truth Differences (PTD), which employs an additional predictor to measure the absolute difference between predicted and true labels, treating larger differences as indicators of higher uncertainty. However, predicting label differences is very difficult, requiring the model to have full discriminative power to estimate the labels of all samples.
  • Ensemble Variance (EV), which trains multiple diverse predictors on different data subsets and uses the variance among their predictions as an uncertainty measure. Although this approach is theoretically sound, it introduces computational overhead because each input must be processed by multiple predictors, conflicting with our efficiency goals.
  • Our Method – Classification-based Entropy (CBE). To overcome the aforementioned limitations, we propose a classification-based entropy approach. Specifically, we transform the regression task into a three-class classification problem and employ a classifier to conduct the classification task, enabling the estimation of predictive uncertainty by computing the entropy of the output probability distribution. This method offers several key advantages: (i) computational efficiency achieved through a single forward pass; (ii) a principled, information-theoretic interpretation of uncertainty; and (iii) reducing the difficulty in uncertainty estimation, as the uncertainty of classification tasks is easier to estimate.
Specifically, alongside the primary regression predictor that outputs continuous sentiment label $\hat{y}_{s}$, we introduce an auxiliary classifier. The continuous sentiment scores are discretized into three categories according to their polarity:
$y_{cls}=\begin{cases}1\;(Negative),&y<0\\ 0\;(Neutral),&y=0\\ 2\;(Positive),&y>0\end{cases}$ (6)
@@
$\begin{split}&\bm{z}_{cls}=Classifier(\bm{h}_{m}),\\ p_{s}=&Softmax(\bm{z}_{cls})=[p_{1},p_{2},p_{3}]\end{split}$ (7)
where ${z}_{cls}$ is a 3-dimensional vector derived from classifier processing, $p_{s}$ denotes the predicted probability distribution over the three sentiment classes.
During inference, uncertainty is quantified by the entropy of the predicted class distribution:
$u_{s}=H(p_{s})=-\sum_{i=1}^{3}p_{s,i}\cdot\log(p_{s,i})$ (8)
where $u_{s}$ denotes the uncertainty score, $H(\cdot)$ is the entropy function, and $p_{s,i}$ is the predicted probability of the $i$-th sentiment class. This entropy-based metric naturally reflects the model’s confidence: a high entropy value indicates greater uncertainty, corresponding to a more uniform distribution over classes, while a low entropy implies strong confidence concentrated on a single class [53].
The model is jointly optimized with a combined loss that integrates both regression and classification objectives:
@@
where $\mathcal{L}_{reg}$ represents the Mean Squared Error (MSE) for regression, $\mathcal{L}_{cls}$ is the Cross-Entropy (CE) loss for classification, and $y_{cls}$ is the discretized ground truth label.
III-D Pipeline of MLLM
We adopt HumanOmni [56] and VideoLLaMA2 [5] as expert oracles to handle high-uncertainty and difficult samples in the MSA task. HumanOmni serves as the core MLLM, whereas VideoLLaMA2 is leveraged for validating the generalization capability of our approach. In the following, we will elaborate on the specific deployment schemes and application scenarios in this study.
The applied HumanOmni and VideoLLaMA2 are both 7-billion-parameter MLLM. These models’ output are formulated as:
$\begin{split}&\mathrm{logits}_{l}=\text{MLLM}(\bm{U}_{t},\bm{U}_{a},\bm{U}_{v}),\\ \text{MLLM}&\in\{\text{HumanOmni},\text{VideoLLaMA2}\}\end{split}$ (10)
where $\bm{U}_{t},\bm{U}_{a},\bm{U}_{v}$ denote the input features from text, audio, and visual modalities, respectively. $\mathrm{logits}_{l}$ is the original probability for samples. Based on the obtained raw probability distribution, we can not only determine the label with the highest prediction likelihood for the corresponding sample, but also derive the probability vector from this distribution to further calculate the uncertainty of each sample:
@@ III-D Pipeline of MLLM
$\begin{split}\hat{y}_{l}=\underset{y_{k}\in y_{l}}{\text{argmax}}\ \text{logits}_{l}(y_{k}),&\\ p_{l}=\mathrm{Softmax}(\mathrm{logits}_{l})=[p_{1},p_{2},&\ldots,p_{n}]\end{split}$ (11)
@@
where $\hat{y}_{l}$ is the predicted sentiment score, $y_{l}$ is the scope of sentiment score, ${p}_{l}$ represents the predicted probability vector which is come from $\mathrm{logits}_{l}$, the definition of $\mathrm{logits}_{l}$ can be seen in Eq. 10. $p_{i}$ is the probability of the $i$-th class, and $n$ is the total number of sentiment classes. We quantify uncertainty by calculating the entropy of this distribution:
$u_{l}=H({p}_{l})=-\sum_{i=1}^{n}p_{l,i}\cdot\log(p_{l,i}),$ (12)
@@
where $u_{l}$ denotes the uncertainty score and $H(\cdot)$ is the entropy function. Higher entropy corresponds to higher uncertainty, providing a principled confidence measure for the model’s predictions. As output consists of multiple tokens, we average the uncertainties of all label-related output tokens to obtain the final uncertainty. The method for determining the uncertainty threshold is explicated in Subsection III-F.
To efficiently adapt HumanOmni and VideoLLaMA2 to our task, we apply LoRA [16], which decomposes the weight update into low-rank matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times d}$, with rank $r$ much smaller than $d$. The adapted weight matrix is given by $W_{new}=W+A\times B$, where $W$ is the original weight matrix and only $A$ and $B$ are trainable during fine-tuning.
III-E Pipeline of U-ACS
Building upon the uncertainty thresholds $\tau_{1}$ and $\tau_{2}$ (see Section III-F), we now present the overall inference workflow, which is shown in Algorithm 1. This workflow employs a dynamic collaboration strategy that sequentially leverages UBM and MLLM, guided by their uncertainty estimates, to efficiently and accurately produce the final prediction. The process proceeds through three stages:
Stage 1: UBM processing. Compute the baseline prediction $\hat{y}_{s}$ and uncertainty $u_{s}$. If the uncertainty is below the adaptive validation-calibrated threshold $\tau_{1}$, accept this prediction directly. Otherwise, proceed to the MLLM Refinement Stage:
$\hat{y}_{1}=\begin{cases}\hat{y}_{s},&u_{s}\leq\tau_{1}\\ \text{Proceed to Stage 2},&u_{s}>\tau_{1}\end{cases}$ (13)
Stage 2: MLLM processing. For samples with higher uncertainty, invoke the MLLM to obtain prediction $\hat{y}_{l}$ along with its uncertainty $u_{l}$. Accept the prediction if the uncertainty is below the second threshold $\tau_{2}$:
$\hat{y}_{2}=\begin{cases}\hat{y}_{l},&u_{l}\leq\tau_{2}\\ \text{Proceed to Stage 3},&u_{l}>\tau_{2}\end{cases}$ (14)
For cases still deemed highly uncertain, apply collaborative strategies based on whether two models agree on sentiment polarity:
  • Agreement (Weighted Averaging): If the predictions of both models share the same sentiment polarity, compute adaptive weights based on uncertainty and combine their outputs:
    $\begin{split}w_{s}&=\frac{1/(u_{s}+\epsilon)}{1/(u_{s}+\epsilon)+1/(u_{l}+\epsilon)},\\ \hat{y}_{3}&=w_{s}\cdot\hat{y}_{s}+(1-w_{s})\cdot\hat{y}_{l}\end{split}$ (15)
  • Disagreement (Cross-Verification): If two models disagree, augment the prompt with previous predictions and uncertainties, and re-predict through the Stage 3.
Stage 3: Collaborative Processsing. Low-confidence samples generated in Stage 2 are sent to Stage 3 for re-prediction, with the prior prediction outputs of Stage 1 and Stage 2 embedded into the prompt as references:
$\begin{split}&\hat{y}_{4}=\text{MLLM}(X,P_{enhanced}),\\ \text{MLLM}&\in\{\text{HumanOmni},\text{VideoLLaMA2}\}\end{split}$ (16)
where $P_{enhanced}$ provides richer context by including UBM’s prediction $\hat{y}_{s}$ and uncertainty $u_{s}$, along with the prediction $\hat{y}_{l}$ and uncertainty $u_{l}$ of the MLLM. $\epsilon$ denotes a small constant parameter with a fixed value of $10^{-12}$.
Finally, we concatenate the output results of three stages to obtain the prediction results for all samples, and calculate the accuracy of our algorithm based on these prediction results:
$\hat{y}=\hat{y}_{1}\|\hat{y}_{2}\|\hat{y}_{3}\|\hat{y}_{4}$ (17)
where $\hat{y}$ denotes the prediction results for all samples, and $\hat{y}_{1}$ represents the prediction result of Stage 1, $\hat{y}_{2}$ and $\hat{y}_{3}$ those of Stage 2, and $\hat{y}_{4}$ that of Stage 3.
III-F Setting Criterion for Uncertainty Threshold
@@
We adopt a principled statistical approach to determine the uncertainty thresholds for both the UBM and the MLLM. This method analyzes the uncertainty distributions on the validation set, providing a theoretically grounded yet computationally efficient way to select thresholds.
@@
The threshold selection procedure is applied consistently to both models. First, we determine the threshold value by calculating the mean of uncertainty values across all samples. The threshold derived from this method has yielded favorable performance in our experiments. To further optimize the results, we finally incorporate the Gaussian fitting strategy and apply this enhanced method to compute the specific thresholds corresponding to each of the two models. For each model, we partition the validation samples into two groups based on whether the model’s predicted polarity matches the ground truth. We define $\mathbb{D}_{same}$ as the set of samples where the prediction and ground truth share the same polarity (both positive or both negative), and $\mathbb{D}_{opposite}$ as the set where they differ in polarity:
@@
where $(u)_{i}$ denotes the uncertainty for sample $i$. Finally, we apply this method with Gaussian fitting to compute the specific thresholds for both models:
@@
Input: Sample $X=(\bm{U}_{t},\bm{U}_{a},\bm{U}_{v})$, UBM $M_{s}$, MLLM $M_{l}$, Uncertainty Thresholds $\tau_{1},\tau_{2}$;
@@
11:   if $\hat{y}_{s}$ and $\hat{y}_{l}$ are the same sentiment polarity then
@@
20:return $\hat{y}=\hat{y}_{1}\|\hat{y}_{2}\|\hat{y}_{3}\|\hat{y}_{4}$
IV Experiments
In this section, we evaluate our strategy on three different multimodal datasets on MSA. We present experiment details and results as follows.
IV-A Datasets
IV-A1 CMU-MOSI
CMU-MOSI (MOSI) [61] contains 2,199 video segments that study the interaction patterns between facial gestures and spoken words for sentiment prediction. The dataset provides fine-grained sentiment annotations on a continuous scale from strongly negative to strongly positive, making it ideal for evaluating both regression and classification capabilities of multimodal models.
IV-A2 CMU-MOSEI
CMU-MOSEI (MOSEI) [62] is a larger multimodal dataset for sentiment analysis and emotion recognition, containing 23,453 annotated video segments from 1,000 distinct speakers across 250 topics, serving as the next generation expansion of MOSI. The dataset’s scale and diversity provide a robust evaluation platform for assessing model generalization across different speakers and topics.
IV-A3 CH-SIMS v1
CH-SIMS v1 (SIMS) [58] is a Chinese multimodal dataset which contains 2,281 refined video segments in the wild with both multimodal and independent unimodal annotation. This dataset enables cross-lingual validation of our approach and demonstrates the framework’s robustness across different languages and cultural contexts.
IV-B Evaluation Metrics
Following standard MSA evaluation protocols, we report comprehensive performance metrics across different granularities:
IV-B1 Classification Metrics